Tickets for a concert were priced at $7 for students and $10 for non-students. There were 2000 tickets sold for a total of $17,339. How may tickets were sold?
Answer by checkley77(12844)

N+S=2,000 OR N=2,000-S
10N+7S=17,339
10(2,000-S)+7S=17,339
20,000-10S+7S=17,339
-3S=17,339-20,000
-3S=-2,661
S=-2,661/-3
S=887 STUDENT TICKETS WERE SOLD.
N=2,000-887=1,113 NON-STUDENT TICKETS WERE SOLD.
PROOF:
10*1,113+7*887=17,339
11,130+6,209=17,339
17,339=17,339